# Are We Asking the Right Questions?: Designing for Community Stakeholders’ Interactions with AI in Policing

### 08 Feb 2024 v1

@@ Abstract.
Research into recidivism risk prediction in the criminal legal system has garnered significant attention from HCI, critical algorithm studies, and the emerging field of human-AI decision-making. This study focuses on algorithmic crime mapping, a prevalent yet underexplored form of algorithmic decision support (ADS) in this context. We conducted experiments and follow-up interviews with 60 participants, including community members, technical experts, and law enforcement agents (LEAs), to explore how lived experiences, technical knowledge, and domain expertise shape interactions with the ADS, impacting human-AI decision-making. Surprisingly, we found that domain experts (LEAs) often exhibited anchoring bias, readily accepting and engaging with the first crime map presented to them. Conversely, community members and technical experts were more inclined to engage with the tool, adjust controls, and generate different maps. Our findings highlight that all three stakeholders were able to provide critical feedback regarding AI design and use - community members questioned the core motivation of the tool, technical experts drew attention to the elastic nature of data science practice, and LEAs suggested redesign pathways such that the tool could complement their domain expertise.
algorithmic crime mapping, human-AI decision-making, problem formulation
@@ 1. Introduction
Predictive policing stands as a significant and contentiously debated topic within the criminal justice system (CJS) today (perry2013predictive, ). To pre-emptively identify where crime is likely to occur and who is likely to commit it (i.e., preventing crime before it happens), police agencies have begun to supplement conventional forecasting and patrolling practices with the use of algorithmic decision-support systems (ADS) (simmons2016quantifying, ). One popular approach, called algorithmic crime mapping (haque2020understanding, ) involves using geographic data to identify hotspots where crime is expected to occur in the future to assist LEAs (Law Enforcement Agents) in proactively and efficiently allocating resources (perry2013predictive, ). Moving to data-driven policing, however, introduces a variety of concerns (borradaile2020whose, ). These systems rely on a large amount of local and geospatial data but fail to account for the trustworthiness or credibility of these different sources of data (clancy2022reconciling, ). Moreover, it is critical to understand the association between crime analysis and how it justifies resource allocation because it is often hard for LEAs to determine how to best incorporate empirical insights within existing work practices in order to improve decision-making (haque2019exploring, ). Moreover, the introduction of ADS in these complex real-world contexts is expected to remove human biases from the decision-making process and make it more objective. However, growing evidence suggests that these systems exacerbate human biases embedded in the historical data (browning2021stop, ; lum2016predict, ; veale2018fairness, ) as well as introduce new biases as a result of the inconsistency between the workers’ decision-making latitude versus the empirical decisions recommended by the ADS (pruss2023ghosting, ; saxena2021framework, ; cheng2022child, ). These issues have underscored the calls for developing human-centered systems (jarrahi2023principles, ; aragon2022human, ; saxena2020human, ) that situate empirical predictions in the context of work and complement workers’ expertise (kawakami2023training, ; saxena2023algorithmic, ), i.e. - allow stakeholders such as law enforcement and crime analysts to derive meaningful information that aligns with their work and informs decision-making on interventions and resource allocation.
@@
However, to ensure that ADS are human-centered, particularly in the public sector, it’s crucial to integrate the perspectives and needs of the stakeholders who are most impacted by the use of such systems, i.e. - community members. Here, SIGCHI researchers have begun collaborating directly with community members affected by ADS (kuo2023understanding, ; stapleton2022has, ; zhang2023deliberating, ; nielsen2023cares, ), in addition to civil servants whose work is impacted by such technology (saxena2023algorithmic, ; kawakami2023training, ; ammitzboll2021street, ). Engaging with a diverse range of community stakeholders enables us to gain a deeper understanding of what it means for ADS used in government agencies to operate in the public interest and further helps us identify ways to enhance the practices of civil servants who are entrusted with serving the community (stapleton2022has, ; wan2023community, ). We contribute to this body of work by introducing an intuitive AI application that community members, technical experts, and law enforcement agents can interact with. The application anchors empirical findings within city neighborhoods familiar to the participants. It served as a boundary object (star1989structure, ) that facilitated detailed and nuanced discussions with them about policing practices in their city and the use of ADS by police departments. Furthermore, it aided in eliciting participants’ feedback on the ethical design and use of such systems.
@@
We conducted a mixed-methods study in a mid-sized Midwestern U.S. city, where we developed an interactive crime-mapping application that facilitated discussions with 60 participants, including community members (n=39), technical experts (n=14), and law enforcement agents (n=7). We developed this application based on prior literature on crime-mapping techniques (chainey_utility_2008, ; johansson2015crime, ; rosser2017predictive, ) where it displays a heatmap resulting from a Kernel Density Estimation (KDE) algorithm (kalinic2018kernel, ) given crime data and specific user inputs. There were two aspects to this study - 1) in a lab activity, participants set alternative parameters for the KDE algorithm and located hotspots on the map, allowing us to assess their interaction and interpretation of algorithmic crime mapping, and 2) through semi-structured interviews conducted immediately after the activity, we gained insights into participants’ perspectives and concerns regarding the use of such tools by police departments as well as their needs that would ensure the ethical design and use of these tools such that they serve the public’s interest. The specific contributions of our paper are –
@@
  • We show that community stakeholders provided critical feedback on various AI life stages: community members shared that the core motivation (i.e., AI problem formulation) was misaligned with the public’s interest, technical experts highlighted the elastic nature of data science practice (i.e., AI development process), and LEAs suggested redesign pathways to complement their domain expertise (i.e., AI interaction and use).
@@
  • We show that intuitive ADS, like crime mapping, can teach community members about the utility, but more importantly, the limitations of AI. They learned that the system was only generating estimates that were subject to change based on their choice of parameters. They were able to contextualize findings based on their local knowledge of the city and deliberate on the significant impact of the underlying data.
@@
  • We found that a target-construct mismatch (kawakami2023training, ) exists even in algorithmic crime mapping, a widely used and intuitive ADS. Law enforcement agents (LEAs) focus on different city landmarks (i.e., the construct) and not the predicted local hotspots (i.e., the target). They prioritize district boundaries and geospatial crime relationships, where the local hotspot only provided a high-level ‘satisfactory outcome’.
@@
  • Consequently, we learned that an ADS that is not worker-centered and failed to consider key attributes valued by LEAs (e.g., relationship between crimes, district boundaries) complicated human-AI decision-making and made it hard to assess the efficacy of human+AI decisions over human-only decisions.
@@
In the following sections, we first discuss related literature on crime-mapping algorithms, involving stakeholders in AI deliberation and design, and human-AI decision-making that have informed this study.
@@ 2.1. AI in Policing and Algorithmic Crime Mapping
Concerning the criminal justice system (CJS), researchers are specifically focused on unraveling stakeholders’ perceptions regarding a myriad of algorithmic systems. This includes understanding judges’ interpretations of diverse risk assessment tools (pruss2023ghosting, ; stevenson2022algorithmic, ), law enforcement’s insights into predictive crime analysis algorithms (veale; haque2019exploring, ), and similar considerations (grgic2018human, ; grgic2016case, ; srivastava2019mathematical, ; 1745-9125.2011.00240, ). A rising number of studies are further attempting to understand how risk assessment algorithms affect different criminal justice outcomes such as pre-trial release, recidivism rates, etc. (green2019principles, ; grgic2019human, ; vaccaro2019effects, ; green2019disparate, ).
@@
Algorithmic crime mapping is the application of contemporary information processing technologies that merge geographic information system (GIS) data, digital maps, and crime data to gain deeper insights into the propagation of criminal activity (haque2020understanding, ). It allows law enforcement organizations to examine and correlate data sources to provide a detailed snapshot of crime episodes and related characteristics within a community or geographic area. (R, ; book12, ). It has already been applied to different types of crime, including drug incidents (12, ), environmental crimes (chainey2013gis, ), burglary (chainey2013gis, ), gang violence (kennedy_known_1997, ), burglary repeat victimization (10.2307/23638645, ), residential burglaries (noauthor_gis_nodate, ), and serial robberies (Hill2002OperationalizingGT, ), among others.
@@
Among various methodologies and tools, spatial clustering and spatiotemporal correlations, spatial ellipses, thematic mapping of geographical regions, grid thematic mapping, and kernel density estimation (KDE) are some more well-known approaches (chainey_utility_2008, ; chainey2013gis, ; 1193063, ; Toole:2011:SCC:1989734.1989742, ). KDE, for example, has lately acquired prominence in terms of practical application and has proven to be quite accurate and predictive (chainey_utility_2008, ). The widely used KDE method involves non-parametric estimation of crime probability density, considering parameters like grid cell size, interpolation methods, and bandwidth for precise hotspot identification (boppuru2020spatio, ). The user’s crucial role involves setting the bandwidth, where a large value may lead to information loss, while a small one emphasizes local data (johansson2015crime, ). This analytical technique, applied to various crime types, transforms low-resolution hotspots into contour line representations with smooth boundaries, increasing generation speed (johansson2015crime, ). Other tools such as non-parametric density estimators for creating maps that are smooth and accurate (chainey2013gis, ; eck_mapping_2005, ), hotspot optimization tool (HOT), made data mining techniques for hotspot mapping easier to implement (Wang:2013:CHM:2584128.2584145, ). Prior research has shown that these tools and algorithms may be used to extract information on criminal organizations from large real-world crime datasets, particularly police-reported crime data, which is very difficult to achieve using standard crime analysis approaches (6425708, ; chainey2013gis, ; 1193063, ).
@@
However, recent SIGCHI research pointed out how such tools and approaches have been linked to several concerns (brayne2017big, ). According to research, racial bias can be found in police reports including racial profiling of vehicles (a, ; doi:10.1080/10439463.2018.1473978, ; doi:10.1080/21565503.2016.1160413, ; doi:10.1111/j.1745-9125.2006.00061.x, ), pedestrian stops (fagan2016stops, ; goel_precinct_2016, ; noauthor_down_nodate, ), traffic fines (dunn2009measuring, ), narcotics enforcement and arrests (lynch2013policing, ; kochel2011effect, ; beckett2006race, ), use of force (buehler2017racial, ; legewie2016racial, ; nix2017bird, ) etc., which can be further amplified if this data has been fed to the prediction tools. The concern is that racially biased police practices will be oriented toward particular places rather than others and knowing that they are in a prediction area may increase police officers’ awareness in ways that accentuate biases such as a minority being subjected to discriminatory police actions (ferguson2012predictive, ). Along with racial bias, privacy concerns have also been raised. (article457, ; 34, ). Because police reports are public information and several police departments offer web-based crime mapping tools on their websites, disseminating spatial crime data might be difficult when the locations of crimes can be linked to specific residences and, thus, specific individuals (34, ). These issues have influenced recent research trends aimed at identifying possible prejudice and unethical implementations.
2.2. Algorithmic Decision-Making in the Public Sector
@@
Algorithmic systems are increasingly being deployed across several public sector organizations such as child welfare (cheng2022child, ; kawakami2022improving, ; saxena2020child, ), public education (mcconvey2023human, ; robertson2021modeling, ), homeless services (moon2024human, ; kuo2023understanding, ; kube2022just, ), criminal justice (stevenson2022algorithmic, ), unemployment services (holten2020shifting, ), and welfare benefits (eubanks2018automating, ) to make high-stakes decisions about citizens’ lives. For instance, algorithms are being used in the criminal justice system to determine sentencing length (grgic2019human, ), allocate resources to neighborhoods (chainey2013examining, ), and predict recidivism (pruss2023ghosting, ). In child welfare, algorithms predict the risk of future maltreatment (cheng2022child, ), determine suitable caregivers (saxena2022unpacking, ), and decide on family services (saxena2023rethinking, ). Public education relies on algorithms for student zoning and performance evaluation (robertson2021modeling, ). Job placement centers use algorithms to profile job seekers and make placement decisions (holten2020shifting, ; ammitzboll2021street, ). Additionally, algorithms establish eligibility criteria and distribute benefits to families in need (eubanks2018automating, ). The deployment and use of algorithms in these sectors have far-reaching consequences, often impacting lives and, in some cases, resulting in life-altering or life-and-death outcomes. As a result, SIGCHI researchers have become increasingly interested in understanding how human-AI decision-making unfolds in these domains (kawakami2022improving, ; saxena2021framework, ), the perspective of frontline workers whose labor is impacted by these technologies (cheng2022child, ; ammitzboll2021street, ), the impact of organizational and policy-related factors (saxena2023algorithmic, ), as well as unpacking implications for the worker-centered design of AI systems in these contexts (saxena2022unpacking, ; kawakami2022care, ).
@@
For instance, researchers have studied how caseworkers in child welfare collaboratively engage with AI systems and how AI-assisted decisions are made within the bounds of organizational and legal constraints (kawakami2022improving, ; saxena2021framework, ; holten2020shifting, ). In a similar vein of work, Kawakami et al. (kawakami2023training, ) develop the notion of critical use that centers the practitioners’ ability to situate AI predictions against the knowledge that is uniquely available to them, i.e. - pre-emptively recognizing that an information asymmetry between the AI and the human is likely to exist and training workers regarding the appropriate reliance on AI systems. Researchers have also explored validity considerations from the social sciences to help improve machine learning problem formulation and assess whether the use of an AI system in a given context is appropriate to begin with (coston2023validity, ; kawakami2023recentering, ). Similarly, Guerdan et al. (guerdan2023ground, ) reviewed several case studies of AI systems in the public sector and developed a causal framework to identify sources of target variable bias and their impact on human-AI decision-making. In addition, researchers have also explored opportunities to support collaboration in human-AI decision-making (buccinca2021trust, ; lai2020chicago, ; green2019disparate, ) by developing explanations (bussone2015role, ), training protocols (kawakami2023training, ), and prescribing away from predicting outcomes (saxena2023algorithmic, ).
@@
In sociotechnical systems, however, AI interfaces rarely contain a clear feedback mechanism that allows users to comprehend the effects of their actions on the system. The growing prevalence of opaque algorithms deployed within public agencies raises concerns about users’ and the community’s education regarding their operation, adoption, and potential biases embedded in them (eslami_understanding_2017, ). We add to the body of knowledge by examining the interaction between a crime-mapping algorithm and users from various backgrounds to identify opportunities and challenges in designing human-centered algorithms that must operate in the public’s interest.
2.3. Community Members’ Participation in AI Deliberation

### 19 Mar 2024 v2

@@ Abstract.
Research into recidivism risk prediction in the criminal justice system has garnered significant attention from HCI, critical algorithm studies, and the emerging field of human-AI decision-making. This study focuses on algorithmic crime mapping, a prevalent yet underexplored form of algorithmic decision support (ADS) in this context. We conducted experiments and follow-up interviews with 60 participants, including community members, technical experts, and law enforcement agents (LEAs), to explore how lived experiences, technical knowledge, and domain expertise shape interactions with the ADS, impacting human-AI decision-making. Surprisingly, we found that domain experts (LEAs) often exhibited anchoring bias, readily accepting and engaging with the first crime map presented to them. Conversely, community members and technical experts were more inclined to engage with the tool, adjust controls, and generate different maps. Our findings highlight that all three stakeholders were able to provide critical feedback regarding AI design and use - community members questioned the core motivation of the tool, technical experts drew attention to the elastic nature of data science practice, and LEAs suggested redesign pathways such that the tool could complement their domain expertise.
algorithmic crime mapping, human-AI decision-making, problem formulation, public sector algorithms
@@ 1. Introduction
Predictive policing stands as a significant and contentiously debated topic within the criminal justice system (CJS) today (Perry, 2013). To pre-emptively identify where crime is likely to occur and who is likely to commit it (i.e., preventing crime before it happens), police agencies have begun to supplement conventional forecasting and patrolling practices with the use of algorithmic decision-support systems (ADS) (Simmons, 2016). One popular approach, called algorithmic crime mapping (Haque et al., 2020) involves using geographic data to identify hotspots where crime is expected to occur in the future to assist LEAs (Law Enforcement Agents) in proactively and efficiently allocating resources (Perry, 2013). Moving to data-driven policing, however, introduces a variety of concerns (Borradaile et al., 2020). These systems rely on a large amount of local and geospatial data but fail to account for the trustworthiness or credibility of these different sources of data (Clancy et al., 2022). Moreover, it is critical to understand the association between crime analysis and how it justifies resource allocation because it is often hard for LEAs to determine how to best incorporate empirical insights within existing work practices in order to improve decision-making (Haque et al., 2019). Moreover, the introduction of ADS in these complex real-world contexts is expected to remove human biases from the decision-making process and make it more objective. However, growing evidence suggests that these systems exacerbate human biases embedded in the historical data (Browning and Arrigo, 2021; Lum and Isaac, 2016; Veale et al., 2018) as well as introduce new biases as a result of the inconsistency between the workers’ decision-making latitude versus the empirical decisions recommended by the ADS (Pruss, 2023; Saxena et al., 2021; Cheng et al., 2022). These issues have underscored the calls for developing human-centered systems (Jarrahi et al., 2023; Aragon et al., 2022; Saxena et al., 2020b) that situate empirical predictions in the context of work and complement workers’ expertise (Kawakami et al., 2023b; Saxena and Guha, 2023), i.e. - allow stakeholders such as law enforcement and crime analysts to derive meaningful information that aligns with their work and informs decision-making on interventions and resource allocation.
@@
However, to ensure that ADS are human-centered, particularly in the public sector, it’s crucial to integrate the perspectives and needs of the stakeholders who are most impacted by the use of such systems, i.e. - community members. Here, SIGCHI researchers have begun collaborating directly with community members affected by ADS (Kuo et al., 2023; Stapleton et al., 2022b; Zhang et al., 2023; Nielsen et al., 2023), in addition to civil servants whose work is impacted by such technology (Saxena and Guha, 2023; Kawakami et al., 2023b; Ammitzbøll Flügge et al., 2021). Engaging with a diverse range of community stakeholders enables us to gain a deeper understanding of what it means for ADS used in government agencies to operate in the public interest and further helps us identify ways to enhance the practices of civil servants who are entrusted with serving the community (Stapleton et al., 2022b; Wan et al., 2023). We contribute to this body of work by introducing an intuitive AI application that community members, technical experts, and law enforcement agents can interact with. The application anchors empirical findings within city neighborhoods familiar to the participants. It served as a boundary object (Star, 1989) that facilitated detailed and nuanced discussions with them about policing practices in their city and the use of ADS by police departments. Furthermore, it aided in eliciting participants’ feedback on the ethical design and use of such systems.
@@
We conducted a mixed-methods study in a mid-sized Midwestern U.S. city, where we developed an interactive crime-mapping application that facilitated discussions with 60 participants, including community members (n=39), technical experts (n=14), and law enforcement agents (n=7). We developed this application based on prior literature on crime-mapping techniques (Chainey et al., 2008; Johansson et al., 2015; Rosser et al., 2017) where it displays a heatmap resulting from a Kernel Density Estimation (KDE) algorithm (Kalinic and Krisp, 2018) given crime data and specific user inputs. There were two aspects to this study - 1) in a lab activity, participants set alternative parameters for the KDE algorithm and located hotspots on the map, allowing us to assess their interaction and interpretation of algorithmic crime mapping, and 2) through semi-structured interviews conducted immediately after the activity, we gained insights into participants’ perspectives and concerns regarding the use of such tools by police departments as well as their needs that would ensure the ethical design and use of these tools such that they serve the public’s interest. In this study, we make the following contributions –
@@
  • We show that community stakeholders are able to provide critical feedback on various AI life stages: community members shared that the core motivation (i.e., AI problem formulation) was misaligned with the public’s interest, technical experts highlighted the elastic nature of data science practice (i.e., AI development process), and LEAs suggested redesign pathways to align AI with practice and complement their domain expertise (i.e., AI interaction and use).
@@
  • We show that intuitive ADS, like crime mapping, can teach community members about the utility, but more importantly, the limitations of AI. They learned that the system only generated estimates that were subject to change based on their choice of parameters. They were able to contextualize findings based on their local knowledge of the city and deliberate on the significant impact of the underlying data.
@@
  • We found that a target-construct mismatch (Kawakami et al., 2023b) exists even in algorithmic crime mapping, a widely used and intuitive ADS. Law enforcement agents (LEAs) focus on different city landmarks (i.e., the construct) and not the predicted local hotspots (i.e., the target). They prioritize district boundaries and geospatial crime relationships, where the local hotspot only provided a high-level ‘satisfactory outcome’.
@@
  • Consequently, we learned that an ADS that is not worker-centered and failed to consider key attributes valued by LEAs (e.g., relationship between crimes, district boundaries) that would augment their practice further complicated human-AI decision-making and made it hard to assess the efficacy of human+AI decisions over human-only decisions.
@@
In the following sections, we first discuss related literature on AI in policing and algorithmic crime mapping, involving stakeholders in AI deliberation and design, and human-AI decision-making that have informed this study.
@@ 2.1. AI in Policing and Algorithmic Crime Mapping
Concerning the criminal justice system (CJS), researchers are specifically focused on unraveling stakeholders’ perceptions regarding a myriad of algorithmic systems. This includes understanding judges’ interpretations of diverse risk assessment tools (Pruss, 2023; Stevenson and Doleac, 2022), law enforcement’s insights into predictive crime analysis algorithms (Veale et al., 2018; Haque et al., 2019), and similar considerations (Grgic-Hlaca et al., 2018, 2016; Srivastava et al., 2019; RATCLIFFE et al., 2011). A rising number of studies are further attempting to understand how risk assessment algorithms affect different criminal justice outcomes such as pre-trial release, recidivism rates, etc. (Green and Chen, 2019b; Grgić-Hlača et al., 2019; Vaccaro and Waldo, 2019; Green and Chen, 2019a).
@@
Algorithmic crime mapping is the application of contemporary information processing technologies that merge geographic information system (GIS) data, digital maps, and crime data to gain deeper insights into the propagation of criminal activity (Haque et al., 2020). It allows law enforcement organizations to examine and correlate data sources to provide a detailed snapshot of crime episodes and related characteristics within a community or geographic area. (Cynthia A. Mamalian and the staff of the Crime Mapping Research Center, [n. d.]; Hirschfield and Bowers, 2001). It has already been applied to different types of crime, including drug incidents (Olligschlaeger, [n. d.]), environmental crimes (Chainey and Ratcliffe, 2013), burglary (Chainey and Ratcliffe, 2013), gang violence (Kennedy et al., 1997), burglary repeat victimization (Johnson et al., 1997), residential burglaries (noa, 2002), and serial robberies (Hill, 2002), among others.
@@
Among various methodologies and tools, spatial clustering and spatiotemporal correlations, spatial ellipses, thematic mapping of geographical regions, grid thematic mapping, and kernel density estimation (KDE) are some more well-known approaches (Chainey et al., 2008; Chainey and Ratcliffe, 2013; Yifei Xue and Brown, 2003; Toole et al., 2011). KDE, for example, has lately acquired prominence in terms of practical application and has proven to be quite accurate and predictive (Chainey et al., 2008). The widely used KDE method involves non-parametric estimation of crime probability density, considering parameters like grid cell size, interpolation methods, and bandwidth for precise hotspot identification (Boppuru and Ramesha, 2020). The user’s crucial role involves setting the bandwidth, where a large value may lead to information loss, while a small one emphasizes local data (Johansson et al., 2015). This analytical technique, applied to various crime types, transforms low-resolution hotspots into contour line representations with smooth boundaries, increasing generation speed (Johansson et al., 2015). Other tools such as non-parametric density estimators for creating maps that are smooth and accurate (Chainey and Ratcliffe, 2013; Eck et al., 2005), hotspot optimization tool (HOT), made data mining techniques for hotspot mapping easier to implement (Wang et al., 2013). Prior research has shown that these tools and algorithms may be used to extract information on criminal organizations from large real-world crime datasets, particularly police-reported crime data, which is very difficult to achieve using standard crime analysis approaches (Tayebi and Glasser, 2012; Chainey and Ratcliffe, 2013; Yifei Xue and Brown, 2003).
@@
However, recent SIGCHI research pointed out how such tools and approaches have been linked to several concerns (Brayne, 2017). According to research, racial bias can be found in police reports including racial profiling of vehicles (Horrace and Rohlin, 2016; Bowling, 2018; Baumgartner et al., 2017; WARREN et al., 2006), pedestrian stops (Fagan et al., 2016; Goel et al., 2016; noa, [n. d.]), traffic fines (Dunn, 2009), narcotics enforcement and arrests (Lynch et al., 2013; Kochel et al., 2011; Beckett et al., 2006), use of force (Buehler, 2017; Legewie, 2016; Nix et al., 2017) etc., which can be further amplified if this data has been fed to the prediction tools. The concern is that racially biased police practices will be oriented toward particular places rather than others and knowing that they are in a prediction area may increase police officers’ awareness in ways that accentuate biases such as a minority being subjected to discriminatory police actions (Ferguson, 2012). Along with racial bias, privacy concerns have also been raised. (Kounadi et al., 2014; Author, 2019a). Because police reports are public information and several police departments offer web-based crime mapping tools on their websites, disseminating spatial crime data might be difficult when the locations of crimes can be linked to specific residences and, thus, specific individuals (Author, 2019a). These issues have influenced recent research trends aimed at identifying possible prejudice and unethical implementations.
2.2. Algorithmic Decision-Support(ADS) Systems in the Public Sector
@@
Algorithmic systems are increasingly being deployed across several public sector organizations such as child welfare (Cheng et al., 2022; Kawakami et al., 2022a; Saxena et al., 2020a), public education (McConvey et al., 2023; Robertson et al., 2021), homeless services (Moon and Guha, 2024; Kuo et al., 2023; Kube et al., 2022), criminal justice (Stevenson and Doleac, 2022), unemployment services (Holten Møller et al., 2020), and welfare benefits (Eubanks, 2018) to make high-stakes decisions about citizens’ lives. For instance, algorithms are being used in the criminal justice system to determine sentencing length (Grgić-Hlača et al., 2019), allocate resources to neighborhoods (Chainey, 2013), and predict recidivism (Pruss, 2023). In child welfare, algorithms predict the risk of future maltreatment (Cheng et al., 2022), determine suitable caregivers (Saxena et al., 2022a), and decide on family services (Saxena et al., 2023). Public education relies on algorithms for student zoning and performance evaluation (Robertson et al., 2021). Job placement centers use algorithms to profile job seekers and make placement decisions (Holten Møller et al., 2020; Ammitzbøll Flügge et al., 2021). Additionally, algorithms establish eligibility criteria and distribute benefits to families in need (Eubanks, 2018). The deployment and use of algorithms in these sectors have far-reaching consequences, often impacting lives and, in some cases, resulting in life-altering or life-and-death outcomes. As a result, SIGCHI researchers have become increasingly interested in understanding how human-AI decision-making unfolds in these domains (Kawakami et al., 2022a; Saxena et al., 2021), the perspective of frontline workers whose labor is impacted by these technologies (Cheng et al., 2022; Ammitzbøll Flügge et al., 2021), the impact of organizational and policy-related factors (Saxena and Guha, 2023), as well as unpacking implications for the worker-centered design of ADS in these contexts (Saxena et al., 2022a; Kawakami et al., 2022b).
@@
For instance, researchers have studied how caseworkers in child welfare collaboratively engage with AI systems and how AI-assisted decisions are made within the bounds of organizational and legal constraints (Kawakami et al., 2022a; Saxena et al., 2021; Holten Møller et al., 2020). In a similar vein of work, Kawakami et al. (Kawakami et al., 2023b) develop the notion of critical use that centers the practitioners’ ability to situate AI predictions against the knowledge that is uniquely available to them, i.e. - pre-emptively recognizing that an information asymmetry between the AI and the human is likely to exist and training workers regarding the appropriate reliance on AI systems. Researchers have also explored validity considerations from the social sciences to help improve machine learning problem formulation and assess whether the use of an AI system in a given context is appropriate to begin with (Coston et al., 2023; Kawakami et al., 2023a). Similarly, Guerdan et al. (Guerdan et al., 2023) reviewed several case studies of AI systems in the public sector and developed a causal framework to identify sources of target variable bias and their impact on human-AI decision-making. In addition, researchers have also explored opportunities to support collaboration in human-AI decision-making (Buçinca et al., 2021; Lai et al., 2020; Green and Chen, 2019a) by developing explanations (Bussone et al., 2015), training protocols (Kawakami et al., 2023b), and prescribing away from predicting outcomes (Saxena and Guha, 2023).
@@
In sociotechnical systems, however, AI interfaces rarely contain a clear feedback mechanism that allows users to comprehend the effects of their actions on the system. The growing prevalence of opaque algorithms deployed within public agencies raises concerns about users’ and the community’s education regarding their operation, adoption, and potential biases embedded in them (Eslami, 2017). We add to the body of knowledge by examining the interaction between a crime-mapping algorithm and users from various backgrounds to identify opportunities and challenges in designing human-centered algorithms that must operate in the public’s interest.
2.3. Community Members’ Participation in the AI Deliberation Process